Grade 9Mechanics

A block of mass 2 kg is joined to a body of mass 1 kg .block is placed on horizontal table and string moves over a Pulley which is the edge of table .mass of 1 kg is suspended by the String. If table is frictionless then acceleration of block and tension in the string are respectively

surendra surendra
7 Years agoGrade 9

1 Answer

Abey
7 Years ago
First consider the 2 blocks as one whole system.
Now a force of 10 N acts on the system because of the hanging 1kg block.
So F=MA, so 10=3×a
a=10/3=3.333
And the tension in the string is 0. As the blocks move freely without friction. 
 
 
 
And the tension